# Sard properties for polynomial maps in infinite dimension

### v2

@@ 4.2. Sard-type theorems and Kolmogorov $n$-width
In this section, we recall the definition of Kolmogorov $n$–width, and we show its role in Sard-type theorems.
Definition 26 (Kolmogorov $n$–width).
Let $X$ be a normed space and $S\subset X$ be a subset. For $n\in\mathbb{N}$, the Kolmogorov $n$–width of $S$ in $X$ is
@@
The asymptotics of $n$–width measures in a quantitative way the compactness of a set, in fact the following holds (see [Pin85, Prop. 1.2]).
@@
$S$ is compact if and only if $S$ is bounded and $\lim_{n\to\infty}\Omega_{n}(S)=0$.
For various properties of the $n$–width we refer the reader to [Pin85]; we recall here those that we will need in the sequel, see [Pin85, Thm. 1.1].
Theorem 28.
@@
As a consequence, in all our results, up to enlarging $S$, one can assume without loss of generality that $S$ is convex and centrally symmetric, without changing its $n$–width.
We establish now a connection between $n$–width of compact sets and locally Lipschitz functions.
@@
Let $\{B_{H}(x_{i},r/2)\}_{i\in I}$ a finite cover of $K$ by balls, with centers $x_{i}\in K$ and radii $r/2>0$. We can assume that $f$ is $L$-Lipschitz on each $B_{H}(x_{i},r)$, for some $L>0$. Since $K$ is compact $\Omega_{n}(K,H)\to 0$ as $n\to\infty$, hence there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ we have $\Omega_{n}(K,H)<r/4$. By definition of $n$–width, for all $n\geq n_{0}$ there exists a $n$-dimensional subspace $E_{n}\subset H$ such that
@@
We can now prove the following Sard-type theorem, corresponding to Theorem A in the Introduction.
@@
Let $q_{\varepsilon}\in(0,q)$. By assumption, $\Omega_{n}(K,H)\leq q^{-n}_{\varepsilon}$ for sufficiently large $n$. We prove that $f$ satisfies the hypothesis of Theorem D. Since $f\in\mathscr{P}^{m}_{d}(H)$ (see Definition 1), the map
@@
is locally Lipschitz. The codomain $Y=\mathbb{R}^{m}\times\mathcal{L}(H,\mathbb{R}^{m})$ is a Banach space equipped with the norm $\|(v,A)\|_{Y}=\|v\|+\|A\|_{\operatorname{op}}$ for $v\in\mathbb{R}^{m}$ and $A\in\mathcal{L}(H,\mathbb{R}^{m})$. Hence we can apply Lemma 30 to get $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ there exists a $n$-dimensional linear subspace $E_{n}\subset H$ such that
@@
As a consequence of the above two estimates, assumption (18) of Theorem D holds, yielding
@@ 4.3. Counterexamples to the Sard theorem in infinite dimension
In this section we provide examples of maps $f\in\mathscr{P}^{m}_{d}(H)$ as in Theorem A for which there exists a compact set $K\subset H$ such that the set $f(\operatorname{Crit}(f)\cap K)$ has not measure zero, however it holds $\Omega_{n}(K,H)\leq cq^{-n}$. This shows the necessity of the quantitative assumption $q>\beta_{0}(m,d)$. As a consequence, we get lower bounds on the semialgebraic constant $\beta_{0}$.
We start with the following example, which is a minor modification of Kupka’s one [Kup65].
@@
where $\phi$ is the Kupka polynomial as in [YC04, Sec. 10.2.3]. Namely $\phi$ has degree $3$, with $\phi(0)=\phi^{\prime}(0)=\phi(1)-1=\phi^{\prime}(1)=0$. If we assume $q^{3}/2<1$ then $f$ is $C^{\infty}$. All critical points have the form
@@
Since $f$ and $\operatorname{Crit}(f)$ satisfy all the hypothesis of Theorem A (provided that $q^{3}/2<1$), we get a consequence that
@@
We need the following classical fact, which follows from [Kuh69]: for any $d\geq 2$ there exists a polynomial $\psi:\mathbb{R}\to\mathbb{R}$ of degree $d$ such that
@@
(See also [Kri84, Thm. 1] which gives a version of this result taking into account also possible multiplicities.) We fix $d\geq 3$ and we consider a polynomial $\psi$ of degree $d$ such that (174) holds. For $q>1$ we consider the map $f_{d}:\ell^{2}\to\mathbb{R}$
@@
Let $\zeta=\max\{|z|\mid\psi^{\prime}(z)=0\}$. Exactly as in Example 32 we get for every $n\in\mathbb{N}$
@@
For any $q\in(1,(d-1)^{1/d})$ the function $f_{d}$ and the set $K=\operatorname{Crit}(f_{d})$ satisfy the hypothesis of Theorem 31, but $f_{d}(\operatorname{Crit}(f_{d}))=[0,1]$. It follows that
@@
Let $f_{d}$ be as in Example 33, and consider $g_{d}:\mathbb{R}^{m-1}\times\ell^{2}\to\mathbb{R}^{m}$ defined as
@@
Taking $E_{n+m-1}=\mathrm{span}\{(e_{i},0),\,(0,e_{j})\mid i=1,\dots,m-1,\,j=1,\dots,n\}\subset\mathbb{R}^{m-1}\times\ell^{2}$, for all $n\in\mathbb{N}$, we can estimate the $n$–width of $\operatorname{Crit}(g_{d})$ as
@@
We observe, provided that $q\in(1,(d-1)^{1/d})$, the function $g_{d}$ satisfies all the hypothesis of Theorem 31, but the set $g_{d}(\operatorname{Crit}(g_{d}))=\mathbb{R}^{m-1}\times[0,1]$ has not measure zero. Hence
@@
We collect the content of Examples 32, 33 and 34 in a unified statement (see Theorem B).
@@
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem A satisfies
@@
We consider $f_{d}:\ell^{2}\to\mathbb{R}$ as in Example 33, and $m\in\mathbb{N}$. We define $h:(\ell^{2})^{m}\to\mathbb{R}^{m}$ as
@@
As in the previous examples, $K$ is compact with exponential $n$–width. By taking finite-dimensional spaces $E_{nm}=\mathrm{span}\{\underbrace{(0,\dots,e_{i},\dots,0)}_{j\text{-th component}}\in(\ell^{2})^{m}\mid i=1,\dots,n,\,j=1,\dots,m\}$, for $n\in\mathbb{N}$ we have
@@
where $\zeta$ is the same constant appearing in Example 33.
@@ 4.4. Sard on linear subspaces
Next, we prove Theorem C, of which we recall the statement.
@@
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem A. Consider the (possibly non-closed) linear subspace
@@
In all previous results, $f\big{(}\mathrm{Crit}(f|_{V})\big{)}$ can be replaced with the smaller set $f(\mathrm{Crit}(f)\cap V)$.
@@
Let $\tilde{H}$ denote the closure of $V$ in $H$. Observe first that
(195) $\Omega_{n}(K,H)=\Omega_{n}(K,\tilde{H}).$
The inequality $\Omega_{n}(K,H)\leq\Omega_{n}(K,\tilde{H})$ is clear since every $n$-dimensional subspace of $\tilde{H}$ is also a subspace of $H$. For the other inequality, since $\tilde{H}$ is closed in $H$, we can write $H=\tilde{H}\oplus\tilde{H}^{\perp}$ and denote by $\pi:H\to\tilde{H}$ the orthogonal projection. Then, using that $K\subset V\subset\tilde{H}$, we obtain
@@
(198) $=\inf_{\begin{subarray}{c}Z\subset\tilde{H}\\ \dim Z\leq n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|$
(199) $=\inf_{\begin{subarray}{c}Z\subset\tilde{H}\\ \dim Z=n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|=\Omega_{n}(K,\tilde{H}).$
We can apply Theorem A to the map $\tilde{f}\in\mathscr{P}_{d}^{m}(\tilde{H})$, defined by $\tilde{f}:=f|_{\tilde{H}}$, with $f\in\mathscr{P}_{d}^{m}(H)$, and the compact set $K\subset\tilde{H}$ obtaining for all $\nu\leq m-1$
@@
Observe now that if $u$ is a critical point for $f|_{V}$, then $u$ is also critical for $\tilde{f}=f|_{\tilde{H}}$. In fact, $D_{u}(f|_{V})=(D_{u}f)|_{V}$ for all $u\in V$. Thus we have $\lambda\in(\mathbb{R}^{m})^{*}\setminus\{0\}$ such that $0\equiv\lambda\circ D_{u}(f|_{V})=(\lambda\circ D_{u}\tilde{f})|_{V}$ and, since $V$ is dense in $\tilde{H}$, it follows that $\lambda\circ D_{u}\tilde{f}\equiv 0$. This means that
(201) $\mathrm{Crit}_{\nu}(f|_{V})\cap K\subseteq\mathrm{Crit}_{\nu}(\tilde{f})\cap K.$
In particular, since on $K\subset V$ we have $\tilde{f}=f|_{V}$ by definition, this implies that
(202) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f|_{V})\cap K\big{)}\bigg{)}\leq\dim_{e}\bigg{(}\tilde{f}\big{(}\operatorname{Crit}_{\nu}(\tilde{f})\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
@@
By Theorem 28 the $n$–width of the compact, convex, centrally symmetric sets $jK^{\prime}$ satisfy
@@
To conclude, observe that any critical point of $f|_{V}$ is also a critical point of $f$ that is also in $V$, thus $f(\mathrm{Crit}(f|_{V}))\subseteq f(\mathrm{Crit}(f)\cap V)$, which yields the last claim in the statement of the theorem. ∎
@@ 4.5. A class of maps with the global Sard property
Using Theorem D, we study a class of maps from a Hilbert space to $\mathbb{R}^{m}$ for which the Sard property holds true. The case $m=1$ is not new, and it was proved in [Yom88], see also [YC04]. We need the following preliminary facts.
@@
From [Kel28, Thm. VI] we get the thesis for $r=1$. Then, we conclude considering the rescaled polynomial $x\to p(rx)$. ∎
@@
where $c_{\alpha}\in\mathbb{R}^{m}$, and the sum is over the multi-indices $\alpha\in\mathbb{N}^{n}$ with $|\alpha|=\sum_{i}\alpha_{i}\leq d$. Since $|\alpha|!c_{\alpha}=\partial_{x^{\alpha}}p|_{x=0}$, we can iterate Markov inequality (for $r=1$, see Lemma 38) and obtain
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1) and its differential is given by
@@
Furthermore, $x\mapsto D_{x}f_{n}\in\mathcal{B}(B_{H}(r);\mathcal{L}(H,\mathbb{R}^{m}))$ for all $n\in\mathbb{N}$, $r>0$. We now prove that it is a Cauchy sequence. Indeed, using Lemma 38, for any $n_{2}\geq n_{1}$
@@
as $n_{2},n_{1}\to\infty$. Therefore, $\{Df_{n}\}_{n\in\mathbb{N}}$ has a limit in $\mathcal{B}(B_{H}(r),\mathcal{L}(H,\mathbb{R}^{m}))$, for all $r>0$. It follows that $f$ is $\mathcal{C}^{1}$ and (215) holds, see e.g. [Lan93, Thm. 9.1].
Thanks to (215), we prove that $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let us consider a sequence $\{x^{j}\}_{j\in\mathbb{N}}$ weakly convergent to $x\in B_{H}(r)$; we prove that $D_{x^{j}}f\to D_{x}f$ as $j\to\infty$. Indeed, given $\varepsilon>0$ we fix $N_{\varepsilon}\in\mathbb{N}$ such that $\sum_{k>N_{\varepsilon}}\sup_{B_{E_{k}}(r)}\|p_{k}\|\leq\varepsilon$ and, by Lemma 38, we get
@@
(221) $\leq\sum_{k=1}^{N_{\varepsilon}}\left\lVert D_{\pi_{k}(x^{j})}p_{k}-D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}+\sum_{k>N_{\varepsilon}}\left\lVert D_{\pi_{k}(x^{j})}p_{k}\right\rVert_{\operatorname{op}}+\sum_{k>N_{\varepsilon}}\left\lVert D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}$
@@
(225) $\leq\frac{\sqrt{m}d^{2}}{r}\sup_{a\in B_{E_{k}}(r)}\|p_{k}(a)-\tilde{p}_{k}(a)\|$ by Lemma 38
@@
(228) $\leq\left(\frac{md^{4}}{r^{2}}\sup_{a\in B_{E_{k}}(3r)}\|p_{k}(a)\|\right)\|z-w\|.$ by Lemma 38
@@
With the following we exhibit a special class of maps, obtained via the construction of Proposition 40, for which the Sard property holds true globally. This is the content of Theorem E in the Introduction, of which we recall the statement here.
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1), and for all $\nu\leq m-1$ and $r>0$ it holds
@@
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
@@
We show first that the map $f$ is well–defined. The upper bound (236), together with Lemma 39 imply that for all $r>0$
@@
We can then apply Proposition 40 obtaining that $f\in\mathscr{P}_{d}^{m}(H)$.
Fix $r>0$. We now prove that $f$ satisfies the hypothesis of Theorem D with $K=B_{H}(r)$ and polynomial approximating maps $f_{n}:E_{n}\to\mathbb{R}^{m}$
@@
(243) $\sup_{K}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}},$

### v3

@@ 4.2. Sard-type theorems and Kolmogorov $n$-width
In this section, we recall the definition of Kolmogorov $n$-width, and we show its role in Sard-type theorems.
Definition 30 (Kolmogorov $n$-width).
Let $X$ be a normed space and $S\subset X$ be a subset. For $n\in\mathbb{N}$, the Kolmogorov $n$-width of $S$ in $X$ is
@@
The asymptotics of $n$-width measures in a quantitative way the compactness of a set, in fact the following holds (see [Pinkus, Prop. 1.2]).
@@
$S$ is precompact if and only if $S$ is bounded and $\lim_{n\to\infty}\Omega_{n}(S)=0$.
For various properties of the $n$-width we refer the reader to [Pinkus]; we recall here those that we will need in the sequel, see [Pinkus, Thm. 1.1].
Proposition 32.
@@
As a consequence, in all our results, up to enlarging $S$, one can assume without loss of generality that $S$ is convex and centrally symmetric, without changing its $n$-width.
We establish now a connection between $n$-width of compact sets and locally Lipschitz functions.
@@
Let $\{B_{H}(x_{i},r/2)\}_{i\in I}$ a finite cover of $K$ by balls, with centers $x_{i}\in K$ and radii $r/2>0$. We can assume that $f$ is $L$-Lipschitz on each $B_{H}(x_{i},r)$, for some $L>0$. Since $K$ is compact $\Omega_{n}(K,H)\to 0$ as $n\to\infty$, hence there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ we have $\Omega_{n}(K,H)<r/4$. By definition of $n$-width, for all $n\geq n_{0}$ there exists a $n$-dimensional subspace $E_{n}\subset H$ such that
@@
We can now prove the following Sard-type theorem, corresponding to Theorem˜A in the Introduction.
@@
Let $q_{\varepsilon}\in(0,q)$. By assumption, $\Omega_{n}(K,H)\leq q^{-n}_{\varepsilon}$ for sufficiently large $n$. We prove that $f$ satisfies the hypothesis of Theorem˜D. Since $f\in\mathscr{P}^{m}_{d}(H)$ (see Definition˜1), the map
@@
is locally Lipschitz. The codomain $Y=\mathbb{R}^{m}\times\mathcal{L}(H,\mathbb{R}^{m})$ is a Banach space equipped with the norm $\|(v,A)\|_{Y}=\|v\|+\|A\|_{\operatorname{op}}$ for $v\in\mathbb{R}^{m}$ and $A\in\mathcal{L}(H,\mathbb{R}^{m})$. Hence we can apply Lemma˜34 to get $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ there exists a $n$-dimensional linear subspace $E_{n}\subset H$ such that
@@
As a consequence of the above two estimates, assumption (19) of Theorem˜D holds, yielding
@@ 4.3. Counterexamples to the Sard theorem in infinite dimension
In this section we provide examples of maps $f\in\mathscr{P}^{m}_{d}(H)$ as in Theorem˜A for which there exists a compact set $K\subset H$ such that the set $f(\operatorname{Crit}(f)\cap K)$ has not measure zero, however it holds $\Omega_{n}(K,H)\leq cq^{-n}$. This shows the necessity of the quantitative assumption $q>\beta_{0}(m,d)$. As a consequence, we get lower bounds on the semialgebraic constant $\beta_{0}$.
We start with the following example, which is a minor modification of Kupka’s one [Kupka].
@@
where $\phi$ is the Kupka polynomial as in [ComteYomdin, Sec. 10.2.3]. Namely $\phi$ has degree $3$, with $\phi(0)=\phi^{\prime}(0)=\phi(1)-1=\phi^{\prime}(1)=0$. If we assume $q^{3}/2<1$ then $f$ is $C^{\infty}$. All critical points have the form
@@
Since $f$ and $\operatorname{Crit}(f)$ satisfy all the hypothesis of Theorem˜A (provided that $q^{3}/2<1$), we get a consequence that
@@
We need the following classical fact, which follows from [realpoly2]: for any $d\geq 2$ there exists a polynomial $\psi:\mathbb{R}\to\mathbb{R}$ of degree $d$ such that
@@
(See also [realpoly, Thm. 1] which gives a version of this result taking into account also possible multiplicities.) We fix $d\geq 3$ and we consider a polynomial $\psi$ of degree $d$ such that (176) holds. For $q>1$ we consider the map $f_{d}:\ell^{2}\to\mathbb{R}$
@@
Let $\zeta=\max\{|z|\mid\psi^{\prime}(z)=0\}$. Exactly as in Example˜36 we get for every $n\in\mathbb{N}$
@@
For any $q\in(1,(d-1)^{1/d})$ the function $f_{d}$ and the set $K=\operatorname{Crit}(f_{d})$ satisfy the hypothesis of Theorem˜35, but $f_{d}(\operatorname{Crit}(f_{d}))=[0,1]$. It follows that
@@
Let $f_{d}$ be as in Example˜37, and consider $g_{d}:\mathbb{R}^{m-1}\times\ell^{2}\to\mathbb{R}^{m}$ defined as
@@
Taking $E_{n+m-1}=\mathrm{span}\{(e_{i},0),\,(0,e_{j})\mid i=1,\dots,m-1,\,j=1,\dots,n\}\subset\mathbb{R}^{m-1}\times\ell^{2}$, for all $n\in\mathbb{N}$, we can estimate the $n$-width of $\operatorname{Crit}(g_{d})$ as
@@
We observe, provided that $q\in(1,(d-1)^{1/d})$, the function $g_{d}$ satisfies all the hypothesis of Theorem˜35, but the set $g_{d}(\operatorname{Crit}(g_{d}))=\mathbb{R}^{m-1}\times[0,1]$ has not measure zero. Hence
@@
We collect the content of Examples˜36, 37 and 38 in a unified statement (see Theorem˜B).
@@
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem˜A satisfies
@@
We consider $f_{d}:\ell^{2}\to\mathbb{R}$ as in Example˜37, and $m\in\mathbb{N}$. We define $h:(\ell^{2})^{m}\to\mathbb{R}^{m}$ as
@@
As in the previous examples, $K$ is compact with exponential $n$-width. By taking finite-dimensional spaces $E_{nm}=\mathrm{span}\{\underbrace{(0,\dots,e_{i},\dots,0)}_{j\text{-th component}}\in(\ell^{2})^{m}\mid i=1,\dots,n,\,j=1,\dots,m\}$, for $n\in\mathbb{N}$ we have
@@
where $\zeta$ is the same constant appearing in Example˜37.
@@ 4.4. Sard on linear subspaces
Next, we prove Theorem˜C. We will need the following lemma.
Lemma 41 (Critical points under restrictions).
Let $(X,\|\cdot\|)$ be a normed space and let $f:X\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map. For any linear subspace $Y\subset X$, consider the restriction $f|_{Y}:Y\to\mathbb{R}^{m}$. Note that the latter is a $\mathcal{C}^{1}$ map in the Fréchet sense from the normed vector space $(Y,\|\cdot\|)$ to $\mathbb{R}^{m}$, with $D_{u}(f|_{Y})=(D_{u}f)|_{Y}$ for all $u\in Y$. Then, for all $\nu=0,\dots,m-1$ it holds
(193) $\mathrm{Crit}_{\nu}(f|_{Y})\supseteq\mathrm{Crit}_{\nu}(f)\cap Y,$
with equality if $Y$ is dense in $X$.
Proof.
The inclusion is proven by noting that, for all $u\in Y$, it holds
(194) $\mathrm{rank}\left(D_{u}(f|_{Y}):Y\to\mathbb{R}^{m}\right)\leq\mathrm{rank}\left(D_{u}f:X\to\mathbb{R}^{m}\right).$
Assume that $Y$ is dense in $X$. Thus if $\lambda\in(\mathbb{R}^{m})^{*}\setminus\{0\}$ is such that $0\equiv\lambda\circ D_{u}(f|_{Y})=(\lambda\circ D_{u}f)|_{Y}$, since $Y$ is dense in $X$, it follows that $\lambda\circ D_{u}f\equiv 0$. This means that for all $u\in Y$
(195) $\mathrm{rank}\left(D_{u}f:X\to\mathbb{R}^{m}\right)\leq\mathrm{rank}\left(D_{u}(f|_{Y}):Y\to\mathbb{R}^{m}\right).$
It follows that $\mathrm{Crit}_{\nu}(f|_{Y})\subseteq\mathrm{Crit}_{\nu}(f)\cap Y$, proving equality of the two sets. ∎
We can now prove Theorem˜C, of which we recall the statement.
@@
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem˜A. Consider the (possibly non-closed) linear subspace
@@
In all previous results, $f\big(\mathrm{Crit}_{\nu}(f|_{V})\big)$ can be replaced with the smaller set $f(\mathrm{Crit}_{\nu}(f)\cap V)$.
@@
Let $X$ denote the closure of $V$ in $H$. Observe first that
(200) $\Omega_{n}(K,H)=\Omega_{n}(K,X).$
The inequality $\Omega_{n}(K,H)\leq\Omega_{n}(K,X)$ is clear since every $n$-dimensional subspace of $X$ is also a subspace of $H$. For the other inequality, since $X$ is closed in $H$, we can write $H=X\oplus X^{\perp}$ and denote by $\pi:H\to X$ the orthogonal projection. Then, using that $K\subset V\subset X$, we obtain
@@
(203) $=\inf_{\begin{subarray}{c}Z\subset X\\ \dim Z\leq n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|$
(204) $=\inf_{\begin{subarray}{c}Z\subset X\\ \dim Z=n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|=\Omega_{n}(K,X).$
We can apply Theorem˜A to the map $\tilde{f}\in\mathscr{P}_{d}^{m}(X)$, defined by $\tilde{f}:=f|_{X}$, with $f\in\mathscr{P}_{d}^{m}(H)$, and the compact set $K\subset X$ obtaining for all $\nu\leq m-1$
@@
By applying Lemma˜41 (with $Y=V$) and noting that, on $K\subset V$, we have $\tilde{f}=f|_{V}$, we obtain:
(206) $\dim_{e}\bigg(f\big(\operatorname{Crit}_{\nu}(f|_{V})\cap K\big)\bigg)=\dim_{e}\bigg(\tilde{f}\big(\operatorname{Crit}_{\nu}(\tilde{f})\cap K\big)\bigg)\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
@@
By Proposition˜32 the $n$-width of the compact, convex, centrally symmetric sets $jK^{\prime}$ satisfy
@@
To conclude, we observe that by Lemma˜41 we have $f(\mathrm{Crit}_{\nu}(f|_{V}))\supseteq f(\mathrm{Crit}_{\nu}(f)\cap V)$, which yields the last claim of the statement of the theorem. ∎
@@ 4.5. A class of maps with the global Sard property
Using Theorem˜D, we study a class of maps from a Hilbert space to $\mathbb{R}^{m}$ for which the Sard property holds true. The case $m=1$ is not new, and it was proved in [YomdinApproxCompl], see also [ComteYomdin]. We need the following preliminary facts.
@@
From [Kellogg1928, Thm. VI] we get the thesis for $r=1$. Then, we conclude considering the rescaled polynomial $x\to p(rx)$. ∎
@@
where $c_{\alpha}\in\mathbb{R}^{m}$, and the sum is over the multi-indices $\alpha\in\mathbb{N}^{n}$ with $|\alpha|=\sum_{i}\alpha_{i}\leq d$. Since $|\alpha|!c_{\alpha}=\partial_{x^{\alpha}}p|_{x=0}$, we can iterate Markov inequality (for $r=1$, see Lemma˜43) and obtain
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition˜1) and its differential is given by
@@
Furthermore, $x\mapsto D_{x}f_{n}\in\mathcal{B}(B_{H}(r);\mathcal{L}(H,\mathbb{R}^{m}))$ for all $n\in\mathbb{N}$, $r>0$. We now prove that it is a Cauchy sequence. Indeed, using Lemma˜43, for any $n_{2}\geq n_{1}$
@@
as $n_{2},n_{1}\to\infty$. Therefore, $\{Df_{n}\}_{n\in\mathbb{N}}$ has a limit in $\mathcal{B}(B_{H}(r),\mathcal{L}(H,\mathbb{R}^{m}))$, for all $r>0$. It follows that $f$ is $\mathcal{C}^{1}$ and (219) holds, see e.g. [LangRealandFunct, Thm. 9.1].
Thanks to (219), we prove that $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let us consider a sequence $\{x^{j}\}_{j\in\mathbb{N}}$ weakly convergent to $x\in B_{H}(r)$; we prove that $D_{x^{j}}f\to D_{x}f$ as $j\to\infty$. Indeed, given $\varepsilon>0$ we fix $N_{\varepsilon}\in\mathbb{N}$ such that $\sum_{k>N_{\varepsilon}}\sup_{B_{E_{k}}(r)}\|p_{k}\|\leq\varepsilon$ and, by Lemma˜43, we get
@@
(228) $\leq\frac{\sqrt{m}d^{2}}{r}\sup_{a\in B_{E_{k}}(r)}\|p_{k}(a)-\tilde{p}_{k}(a)\|$ by Lemma˜43
@@
(231) $\leq\left(\frac{md^{4}}{r^{2}}\sup_{a\in B_{E_{k}}(3r)}\|p_{k}(a)\|\right)\|z-w\|.$ by Lemma˜43
@@
With the following we exhibit a special class of maps, obtained via the construction of Proposition˜45, for which the Sard property holds true globally. This is the content of Theorem˜E in the Introduction, of which we recall the statement here.
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition˜1), and for all $\nu\leq m-1$ and $r>0$ it holds
@@
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem˜D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
@@
We show first that the map $f$ is well–defined. The upper bound (239), together with Lemma˜44 imply that for all $r>0$
@@
We can then apply Proposition˜45 obtaining that $f\in\mathscr{P}_{d}^{m}(H)$.
Fix $r>0$. We now prove that $f$ satisfies the hypothesis of Theorem˜D with $B=B_{H}(r)$ and polynomial approximating maps $f_{n}:E_{n}\to\mathbb{R}^{m}$
@@
(246) $\sup_{x\in B}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}},$